# 2D Entanglement from a Single Motzkin Spaghetto

### v1

@@ Abstract
While highly entangled ground states of gapless local Hamiltonians have been known to exist in one dimension, their 2D counterparts were only recently found, with rather sophisticated interactions involving at least 4 neighboring degrees of freedom. In this letter, a 2D lattice system interacting only between two nearest neighboring sites is proposed, which has a unique ground state with volume scaling of entanglement entropy. The system undergoes an entanglement phase transition with enhanced entanglement entropy at the critical point compared to the previous construction from coupling orthogonal arrays of one-dimensional chains, but the spectral gap in the highly entangled phase decays faster. A tensor network representation is composed of tensors with lower rank, and has a similar global geometry.
@@
Introduction—Entanglement entropy (EE) is a measure of how entangled a quantum state is, or how different it is from a classical state. In that sense, the ground state (GS) of a locally interacting quantum many-body system separated by a spectral gap from the first excited state is not so different from a classical state, at least compared to generic eigenstates in the Hilbert space [1], as the EE of a subsystem grows with the size of the boundary instead of the bulk. This is known as the area law of EE [2], and is expected to hold in any spatial dimension. It is intuitive because massive quasiparticles propagate at a finite speed, resulting in exponentially decaying correlations. So only degrees of freedom within a finite range are effectively entangled across a boundary and contribute to the bipartite EE. The idea has been made rigorous into a proof for 1D systems using Lieb-Robinson bounds [3]. However, in 2D progress towards establishing a theorem has only been made for certain class of Hamiltonians known as frustration free [4], to which the model proposed in this letter belongs.
@@
What the area law does not tell us is how entangled a GS can be when the spectral gap closes, as happens in quantum criticality. There, the EE typically scales with an extra factor logarithmic in the subsystem size beyond the area law. This has been demonstrated with conformal field theories for (1+1)D critical points [5, 6], and with the Widom conjecture for gapless free fermions with a Fermi sea in arbitrary dimensions [7]. That was perhaps not so surprising as correlation is known to have a power law decay at critical points. But what was a little striking was that when the spectral gap decays even faster with the system size, for example at an exponential rate as opposed to the polynomial decay for critical systems, the EE can actually have a volume scaling. This was first exhibited in an inhomogeneously interacting XX spin chain [8, 9], where the spins can be approximated to form Bell pairs symmetrically across the center in the GS, leading to maximal EE. The model was hence referred to as the rainbow chain.
@@
Similar examples of highly entangle GSs with more exact descriptions were made possible with frustration free Hamiltonians [10], that are also translationally invariant (broken only by the finite boundary). For those Hamiltonians, the GS is simultaneously the lowest energy eigenstate of each local interaction operator, such that a lot can be known analytically about the GS by diagonalizing a few-body Hamiltonian. By mapping spin-1 configurations to discrete random walks, the GS of the critical Hamiltonian in Ref. [10] can be given as a superposition of Motzkin paths, those that start and end at hight zero, without ever passing below it in the middle. Using combinatorial results, Bravyi et al. showed that the GS of the Motzkin chain has the same scaling behavior for EE, correlation and spectral gap as typical critical systems. Movassagh and Shor enlarged the local Hilbert space of the Motzkin chain with a color degree of freedom, which has a power law scaling of EE that outgrows the spin sector [11]. These constructions were later generalized to spin-$\frac{1}{2}$ systems, with smaller local degrees of freedom, but three-body interaction involving next-nearest neighboring sites [12, 13], giving its name Fredkin chain. To actually realize volume scaling in the GS, these Hamiltonians have to be deformed in a frustration free way, such that the GS superposition favors higher paths which allow more Bell pairs across the bipartition boundary [14, 15, 16]. As the deformation parameter is tuned, the GS undergoes an entanglement phase transition from area law to volume scaling.
@@
While the six- or 19-vertex and lozenge tiling realizations appear to be the natural generalization of the highly entangled Motzkin (resp. Fredkin) chain to 2D, the support of the local Hamiltonian operators grows from two (resp. three) in 1D to four (resp. six). So while the paradigm in principle works for higher dimensions, the interactions become increasingly less realistic. An alternative loop-soup construction was proposed in Ref. [18], where the GS is a superposition of random loop configurations on a 2D lattice, and on each loop lives a periodic Motzkin chain [23]. However, due to the randomness of the loop configurations, it is not clear how the EE for any bipartition would actually scale with the subsystem size. Moreover, a naive generalization of the Motzkin Hamiltonian to periodic boundary conditions leaves the GS with exponential degeneracy, making a combination with exotic EE scaling trivially exist. To avoid both of those complications, we focus on a 2D lattice filled by winding one open 1D chain. The interaction is turned on for bonds along the chain, so it remains two- or three-body. Since the location of the entangling chain is fixed, we are able to argue from the 1D results that the EE scaling is asymptotically the same as the coupled-chain construction, except at the critical point, and the entanglement is isotropic when the subsystems are partitioned in different directions. Furthermore, a tensor network (TN) representation of the new GS using the previously obtained TNs for the 1D chains [24, 25] has the same global geometry as the TN for the GS of the six-vertex coupled Fredkin chains [26].
@@
Motzkin chains—The $q$-deformed bicolor Motzkin chain [14] has a five-dimensional local Hilbert space, with spin up and down, both in either red (r) or blue (b) color, and a colorless spin zero. The Hamiltonian is expressed in terms of projectors $P_{j}^{c}=|L^{c}\rangle_{j,j+1}\langle L^{c}|_{j,j+1}+|R^{c}\rangle_{j,j+1}\langle R^{c}|_{j,j+1}+|F^{c}\rangle_{j,j+1}\langle F^{c}|_{j,j+1}$, where
@@
Each of the projection operator has a local lowest energy eigenstate corresponding to the direction orthogonal to the above three vectors. So the GS of a Hamiltonian consisting of such projectors acting on all neighboring pairs of sites in the chain is a weighted superposition of random walks if the up, down and 0 spins are respectively mapped to up, down and flat moves. In addition to the color matching encoded in $|F^{c}\rangle_{j,j+1}$ between up and down moves when they are pair created from spin zeros, an energy penalty needs to be introduced by the projector $C_{j}^{c_{1},c_{2}}=|\uparrow^{c_{1}}_{j}\downarrow^{c_{2}}_{j+1}\rangle\langle\uparrow^{c_{1}}_{j}\downarrow^{c_{2}}_{j+1}|$ for $c_{1}\neq c_{2}$ in order to enforce color matching between nearest up-down pairs in the GS configurations. The remaining GS degeneracy can be lifted by fixing the boundary conditions with the Hamiltonian $H_{\partial}=\sum_{c=\mathrm{r,b}}|\downarrow^{c_{1}}\rangle\langle\downarrow^{c_{1}}|_{1}+|\uparrow^{c_{1}}\rangle\langle\uparrow^{c_{1}}|_{2N}$. In summary, the Motzkin Hamiltonian is given by
@@
2D lattice—Since the 2D model is constructed by winding the 1D chain, we only have to specify the location of the 1D chain on a 2D grid. This is shown in Fig. 1. The spin degrees of freedom reside only on the edges of the square lattice covered by the winding 1D chain, and interactions are only turned on between adjacent spins on the chain. Depending on the location of the spin in the square lattice, it can be either between east-west (resp. north-south) neighbors for the upper and lower (resp. left and right) quadrants, or neighbors around a corner near the two diagonals.
Figure 1: A twisted Motzkin chain of length $2N=L(L+2)$ occupying an $L\times L$ square lattice, shown here for $L=8$. The spin degrees of freedoms live on the edges of the square lattice, as are labeled in the figure. Four symmetric bipartition cuts are shown in different colors.
To study the isotropic EE, four symmetric bipartition cuts are shown in Fig. 1. They separate the spins into two subsystems consisting of increasingly larger segments of size $4n-a$, for $n=1,2,\cdots,L/2$, with $a=0,1,2,3$ corresponding respectively to the yellow, blue, green and red cut. Starting from the center, segments within one subsystem in the 2D lattice on a given side of the cut alternate between the first and second halves of the 1D chain when it is stretched straight. Obviously the scaling behavior of bipartite EE for any of the cuts will not differ qualitatively, so from now on we focus on the cut along the main diagonal with $a=0$, which results in a bipartition shown in Fig. 2.
@@
Figure 2: The two subsystems resulting from the yellow cut in Fig. 1. The northeast subsystem is represented with the thickened black lines, whereas the southwest is marked by thin gray lines. The configuration depicted corresponds to the probabilistically dominant Motzkin path in the GS superposition for the $q>1$ phase, to illustrate the entanglement across subsystems. Typical height configuration for at the critical point is depicted in blue.
@@
Scaling of EE—As analyzed in Ref. [14], the EE decomposes into a leading contribution from the color degree of freedom, and a subleading contribution from the spin sector. Only up-down spin pairs belonging to different subsystems account for the former. In the Motzkin path picture, this means that only the net height change of a subsystem, or the surplus of up spins matters. For the 1D system studied there, the leading contribution to the EE is therefore proportional to the height at the center of the Motzkin path, giving the three distinct scaling behaviors in the two different phases and the critical point.
@@
Here, the situation is a little more complicated, as segments belonging to two subsystems alternate along the chain. But we can get lower or upper bounds for the two phases with the following argument. For $q>1$, the net height change within each segment scales linearly with its length, at least for segments long enough which are the overwhelming majority in the asymptotic limit. If there is a dip in the middle of the segment, it would only increase the entanglement between the two subsystems, as the immediate neighboring segment on the left belongs to the other subsystem. For down spin to actually form up-down pairs with up spins from the next neighboring segment to the left, coming from the same subsystem, the probability would be of the next subleading order. Therefore, assuming the height to monotonically increase for the first half of the chain gives a lower bound on the EE, which is already growing linearly following the proof for the 1D case [14].
@@
For $q<1$, the net height fluctuation stays at order $\mathcal{O}(1)$ anywhere along the chain. In the best case scenario for accumulating EE, the height changes are positive for all the segments in one subsystem, and negative in the other. Since there are in total $L/2$ segments of each, the EE scales at most with the linear size of the lattice, which means the area law.
The $q=1$ case calls for more careful analysis, where the hard wall constraint from below determines the shape of typical Motzkin paths. Recall that the most probable height of a Motzkin path that starts from zero height scales as $h\approx\sqrt{x/3}$ after $x$ steps [10]. Following a completely analogous enumeration of discrete paths using the reflection method, and then taking the asymptotic limit, it can be shown that a Motzkin path that starts at height $h_{0}$ typically ends at height $h_{1}\approx\sqrt{(x-h_{0}^{2})/3}$ after $x$ steps. Iteratively applying this formula shows the height change within each segment scales as $\sqrt{L}$. Since there are order $L$ sufficiently long segments, the total EE scales as $L^{3/2}$, different from the $L\log L$ scaling of the couple chain constructions [20, 21].
@@
The subleading contribution to the EE is from the fluctuation in the height of Motzkin paths, which is also the sole source of entanglement for a monochromatic version of the model. A Schmidt decomposition of the reduced density matrix shows that the two subsystems only depend on each other via the total spin within one subsystem, such that the total spin of the whole chain is zero. Moving from the 1D system to our 2D construction, nothing changes except the subsystem is distributed into disconnected segments. So dimensionality of the reduced density matrix is still the number of possible values of the total spin. However, in the $q>1$ (resp. $q<1$) phase, the probability of the total spin, or height change decays exponentially with the deviation of the area underneath a Motzkin path from the most probable one that is of highest height (resp. completely flat) [14, 15]. So only paths with a height change at finite distance away from the typical one effectively matters, given a bounded EE independent of system size for both phases. When $q=1$, there is more fluctuation. Although entropic repulsion pushes the average height to order $\sqrt{N}$ above the hard wall at zero height, the variance of the Gaussian free field remains to scale linearly with the total length of the segments, making the EE scale as $\log N$ [10]. Note that in a monochromatic version of the 2D model, all three scalings are sub-area law, which is unique to our construction of a 2D lattice by winding a 1D chain. The EE scalings for the two different 2D models are compared in Table 1 for different phases.
@@
Spectral gap—Unlike the scaling of EE discussed above, which requires somewhat different analyses than the 1D system due to the more complicated bipartition, upper-bounds on the scaling of the spectral gap can be straightforwardly deduced from the 1D scaling simply by substituting the linear system size with $N=L(L+2)$. The results are summarized in Table 2. This should be compared with the EE scaling in Table 1. In particular, the $q>1$ phase of the colored model shows that the same EE scaling does not imply the same scaling of spectral gap, and vice versa as shown by the critical point of the colored model and the $q>1$ phase of the colorless one.

### v2

# 2D Entanglement from a Twisted Motzkin Spaghetto

@@ Abstract
While highly entangled ground states of gapless local Hamiltonians have been known to exist in one dimension, their 2D counterparts were only recently found, with rather sophisticated interactions involving at least 4 neighboring degrees of freedom. Here, we show that similar 2D entanglement properties can be realized by twisting a single 1D chain embedded on a 2D lattice. The interaction to generate such 2D entanglement is much simpler than the previous 2D construction by coupling orthogonal arrays of 1D highly entangled chains. The new construction exhibits an entanglement phase transition with enhanced entanglement entropy at the critical point compared to the previous bona fide 2D models, but the spectral gap in the highly entangled phase decays faster. The tensor network representation of the new ground state is composed of tensors with lower rank, showing a similar global geometry as those of the old ones.
I Introduction
@@
Entanglement entropy (EE) is a measure of how entangled a quantum state is, or how different it is from a classical state. In that sense, the ground state (GS) of a locally interacting quantum many-body system separated by a spectral gap from the first excited state is not so different from a classical state, at least compared to generic eigenstates in the Hilbert space [1], as the EE of a subsystem grows with the size of the boundary instead of the bulk. This is known as the area law of EE [2], and is expected to hold in any spatial dimension. It is intuitive because massive quasiparticles propagate at a finite speed, resulting in exponentially decaying correlations. So only degrees of freedom within a finite range are effectively entangled across a boundary and contribute to the bipartite EE. The idea has been made rigorous into a proof for 1D systems using Lieb-Robinson bounds [3]. However, in 2D progress towards establishing a theorem has only been made for a certain class of Hamiltonians known as frustration free [4], to which the model proposed in this letter belongs.
@@
What the area law does not tell us is how entangled a GS can be when the spectral gap closes, as happens in quantum criticality. There, the EE typically scales with an extra factor logarithmic in the subsystem size beyond the area law. This has been demonstrated with conformal field theories for (1+1)D critical points [5, 6], and with the Widom conjecture for gapless free fermions with a Fermi sea in arbitrary dimensions [7]. It is perhaps not so surprising as the correlation length typically scales linearly with the system size at critical points. But what was a little striking was that when the spectral gap decays even faster with the system size, for example at an exponential rate as opposed to the polynomial decay for critical systems, the EE can actually have a volume scaling. This was first exhibited in an inhomogeneously interacting XX spin chain [8, 9, 10], where the spins can be approximated to form Bell pairs symmetrically across the center in the GS, leading to maximal EE. The model was hence referred to as the rainbow chain.
@@
Similar examples of highly entangled GSs with more exact descriptions were made possible with frustration free Hamiltonians [11], that are also translationally invariant (broken only by the finite boundary). For those Hamiltonians, the GS is simultaneously the lowest energy eigenstate of each local interaction operator, such that a lot can be known analytically about the GS by diagonalizing a few-body Hamiltonian. By mapping spin-1 configurations to discrete random walks, the GS of the critical Hamiltonian in Ref. [11] can be given as a superposition of Motzkin paths, those that start and end at height zero, without ever passing below it in the middle. Using combinatorial results, Bravyi et al. showed that the GS of the Motzkin chain has the same scaling behavior for EE, correlation and spectral gap as typical critical systems. Movassagh and Shor enlarged the local Hilbert space of the Motzkin chain with a color degree of freedom, which has a power law scaling of EE that outgrows the spin sector [12]. These constructions were later generalized to spin-$\frac{1}{2}$ systems, with smaller local degrees of freedom, but three-body interaction involving next-nearest neighboring sites [13, 14], giving its name Fredkin chain. To actually realize volume scaling in the GS, these Hamiltonians have to be deformed in a frustration free way, such that the GS superposition favors higher paths which allow more Bell pairs across the bipartition boundary [15, 16, 17]. As the deformation parameter is tuned, the GS undergoes an entanglement phase transition from area law to volume scaling.
@@
While the six- or 19-vertex and lozenge tiling realizations appear to be the natural generalization of the highly entangled Motzkin (resp. Fredkin) chain to 2D, the support of the local Hamiltonian operators grows from two (resp. three) in 1D to four (resp. six). So while the paradigm in principle works for higher dimensions, the interactions become increasingly less realistic. There has also been a proposal of sequentially generable 2D coupled GS that is highly entangled in only one direction with local Hamiltonian operators having even larger support of eight [24]. An alternative loop-soup construction was proposed in Ref. [19], where the GS is a superposition of random loop configurations on a 2D lattice, and on each loop lives a periodic Motzkin chain [25]. However, due to the randomness of the loop configurations, it is not clear how the EE for any bipartition would actually scale with the subsystem size. Moreover, a naive generalization of the Motzkin Hamiltonian to periodic boundary conditions leaves the GS with exponential degeneracy, making a combination with exotic EE scaling trivially exist. To avoid both of those complications, we focus on a twisting open 1D chain from the center to fill a 2D lattice, such that entanglement in both directions of the 2D lattice is isotropic. At the same time, the interaction is turned on only for 2D lattice bonds that reside along the twisted chain, so it remains two or three nearest neighbors. Since the location of the entangling chain is fixed, we are able to argue from the 1D results that the EE scaling is asymptotically the same as the coupled-chain construction, except at the critical point.
The EE scaling of a quantum state is closely related to its tensor network (TN) representation. A TN with hierarchical structure for the Motzkin GS with logarithm violation of the area law proposed in Ref. [26] has recently been shown to be equivalent to a generalized multiscale entanglement renormalization ansatz (MERA) [27]. The equivalence was established by decomposing a rank-5 tensor, the minimal unit cell of the TN, into a product of rank-3 tensors and recombining them with their neighbors. Here, we perform a similar reshaping of the TN for the volume entangled Fredkin chain inspired by the recently discovered holographic TN for the double-scaled SYK model [28]. By considering the conservation of arrow flows, we further optimize the tensors to a smaller set of configurations different on even and odd coordinates in the 2D bulk of the holographic TN. The new TN is then used to describe the GS of the current 2D model, showing global similarity with the previously obtained TN for the GS of the six-vertex coupled Fredkin chains [29].
The rest of the article is organized as follows. Sec. II defines the model with highly entangled GS in two directions by filling a 2D square lattice with a Motzkin chain twisted from the center. Sec. III analyzes the asymptotic scaling of the EE by mapping bipartitions of the 2D system into disjoint segments along the 1D chain. The result together with the scaling behavior of spectral gap converted from the 1D results for different phases and the critical points are compared with those for the previous coupled-chain construction. Sec. IV highlights the induced antiferromagnetic order in the spin sector along the radial direction of the 2D lattice. Sec. V introduces a new TN representation of the highly entangled 1D chains, the 2D twisted version of which is compared with the genuine 2D TN with similar global structure. Finally, a conclusion is given in Sec. VI with outlook on higher dimensional generalizations and EE scaling laws for gapless GSs.
II Twisted Motzkin chain
@@
The $q$-deformed bicolor Motzkin chain [15] has a five-dimensional local Hilbert space, with spin up and down, both in either red (r) or blue (b) color, and a colorless spin zero. The Hamiltonian is expressed in terms of projectors $P_{j}^{c}=|L^{c}\rangle_{j,j+1}\langle L^{c}|_{j,j+1}+|R^{c}\rangle_{j,j+1}\langle R^{c}|_{j,j+1}+|F^{c}\rangle_{j,j+1}\langle F^{c}|_{j,j+1}$, where
@@
Each of the projection operators has a local lowest energy eigenstate corresponding to the direction orthogonal to the above three vectors. So the GS of a Hamiltonian consisting of such projectors acting on all neighboring pairs of sites in the chain is a weighted superposition of random walks if the up, down and 0 spins are respectively mapped to up, down and flat moves. In addition to the color matching encoded in $|F^{c}\rangle_{j,j+1}$ between up and down moves when they are pair created from spin zeros, an energy penalty needs to be introduced by the projector $C_{j}^{c_{1},c_{2}}=|\uparrow^{c_{1}}_{j}\downarrow^{c_{2}}_{j+1}\rangle\langle\uparrow^{c_{1}}_{j}\downarrow^{c_{2}}_{j+1}|$ for $c_{1}\neq c_{2}$ in order to enforce color matching between nearest up-down pairs in the GS configurations. The remaining GS degeneracy can be lifted by fixing the boundary conditions with the Hamiltonian $H_{\partial}=\sum_{c=\mathrm{r,b}}|\downarrow^{c_{1}}\rangle\langle\downarrow^{c_{1}}|_{1}+|\uparrow^{c_{1}}\rangle\langle\uparrow^{c_{1}}|_{2N}$. In summary, the Motzkin Hamiltonian is given by
@@
Since the 2D model is constructed by twisting a 1D chain quadratically as long, we only need to specify the bonds on the 2D lattice occupied by degrees of freedom along the 1D chain. This is shown in Fig. 1 (a). The colored spins reside only on the edges of the square lattice covered by the twisted 1D chain, and interactions are only turned on between adjacent spins on the chain. Depending on the location of the spin in the square lattice, it can be either between east-west (resp. north-south) neighbors for the upper and lower (resp. left and right) quadrants. Near the two diagonal lines, where the winding 1D chain turns, the interaction is between neighbors around a corner.
@@
Figure 1: (a) A twisted Motzkin chain of length $2N=L(L+2)$ embedded in an $L\times L$ square lattice, shown here for $L=8$. The spin degrees of freedoms live on the edges of the square lattice, as are numbered in the figure. Four symmetric bipartition cuts are shown in different colors. (b) The two subsystems resulting from the yellow cut in (a). The northeast subsystem is represented with the thickened black lines, whereas the southwest is marked by thin gray lines. The configuration depicted corresponds to the probabilistically dominant Motzkin path in the GS superposition for the $q>1$ phase, to illustrate the entanglement across subsystems. Typical height configuration at the critical point is depicted in blue.
To study the isotropic EE, four symmetric bipartition cuts are shown in Fig. 1 (a). They separate the spins into two subsystems consisting of increasingly larger segments of size 4na4n-a, for $n=1,2,\cdots,L/2+1$, with $a=0,1,2,3$ corresponding respectively to the yellow, blue, green and red cut. Starting from the center, segments within one subsystem in the 2D lattice on a given side of the cut alternate between the first and second halves of the 1D chain when it is stretched straight. Obviously the scaling behavior of bipartite EE for any of the cuts will not differ qualitatively, so from now on we focus on the cut along the main diagonal with $a=0$, which results in a bipartition shown in Fig. 1 (b).
III Scaling of EE and spectral gap
@@
As analyzed in Ref. [15], the EE of the Motzkin GS decomposes into a leading contribution from the color degree of freedom, and a subleading contribution from the spin sector. Only up-down spin pairs belonging to different subsystems account for the former. In the Motzkin path picture, this means that only the net height change of a subsystem, or the surplus of up spins matters. For the 1D system studied there, the leading contribution to the EE is therefore proportional to the height at the center of the Motzkin path, giving the three distinct scaling behaviors in the two different phases and the critical point. For the critical point at $q=1$, the EE of a single connected segment in the middle of the chain has also been obtained from combinatorial analyses for the colorless [30] and colored Motzkin chains [31].
@@
Here, the situation is a bit more complicated, as segments belonging to two subsystems alternate along the chain. But we can get lower or upper bounds for the two phases with the following argument. For $q>1$, the net height change within each segment scales linearly with its length, at least for segments long enough which are the overwhelming majority in the asymptotic limit. If there is a dip in the middle of the segment, it would only increase the entanglement between the two subsystems, as the immediate neighboring segment on the left belongs to the other subsystem. For down spins to actually form up-down pairs with up spins from the next neighboring segment to the left, coming from the same subsystem, the probability would be of the next subleading order. Therefore, assuming the height to monotonically increase for the first half of the chain gives a lower bound on the EE, which is already growing linearly following the proof for the 1D case [15].
@@
For $q<1$, the net height fluctuation stays at order $\mathcal{O}(1)$ anywhere along the chain. In the best case scenario for accumulating EE, the height changes are positive for all the segments in one subsystem, and negative in the other. Since there are in total $L/2$ segments of each, the EE scales at most with the linear size of the lattice, which means the area law for the 2D system.
At $q=1$, the net height change for a segment of size $4n$ typically scales as $\sqrt{4n}$. This can be understood as the standard deviation of a random walk of $4n$ steps, and is qualitatively unaffected by the hard wall constraint from below. So the total number of up-down spins to be color coordinated across two subsystems can be estimated as $\sum_{n=1}^{L/2}\sqrt{4n}$, which scales as $L^{3/2}$. The difference from the $L\log L$ scaling of the couple chain constructions [21, 22] can be understood from the enhanced fluctuation of height in 1D interaction compared to 2D.
@@
The subleading contribution to the EE is from the fluctuation in the height of Motzkin paths, which is also the sole source of entanglement for a monochromatic version of the model. A Schmidt decomposition of the reduced density matrix shows that the two subsystems only depend on each other via the total spin within one subsystem, such that the total spin of the whole chain is zero. Moving from the 1D system to our 2D construction, nothing changes except the subsystem is distributed into disconnected segments. So dimensionality of the reduced density matrix is still the number of possible values of the total spin. However, in the $q>1$ (resp. $q<1$) phase, the probability of the total spin, or height change decays exponentially with the deviation of the area underneath a Motzkin path from the most probable one that is of highest height (resp. completely flat) [15, 16]. So only paths with a height change at finite distance away from the typical one effectively matters, given a bounded EE independent of system size for both phases. When $q=1$, there is more fluctuation. Although entropic repulsion pushes the average height to order $\sqrt{N}$ above the hard wall at zero height, the variance of the Gaussian free field remains to scale linearly with the total length of the segments, making the EE scale as $\log N$ [11]. Note that in a monochromatic version of the 2D model, all three scalings are sub-area law, which is unique to our construction of filling the 2D lattice with a twisted 1D chain. The EE scalings for the two different 2D models are compared in Table 1 for different phases.
@@
Unlike the scaling of EE discussed above, which requires somewhat different analyses than the 1D system due to the more complicated bipartition, upper-bounds on the scaling of the spectral gap can be straightforwardly deduced from the 1D scaling simply by substituting the linear system size with $N=L(L+2)$. The results are summarized in Table 1. This should be compared with the EE scaling in Table 1. In particular, the $q>1$ phase of the colored model shows that the same EE scaling does not imply the same scaling of spectral gap, and vice versa as shown by the critical point of the colored model and the $q>1$ phase of the colorless one.
IV Antiferromagnetic order
Figure 2: Order parameter $\langle S^{z}\rangle$ for (a) $q>1$, and (b) $q=1$. Warm color and cold colors denote respective spin up and down, and darker colors correspond to larger magnitude. The neutral muted shade represents 2D lattice sites not occupied by degrees of freedom of the 1D chain.
Although the twisted 1D chain fills the entire 2D lattice in our construction, one might argue that the system is not a genuine interacting in 2D, as there are no couplings between adjacent segments of the 1D chain within each quadrant of the lattice. The caveat in that argument is that the coupling within different quadrants are along different directions, vertical in the east and west quadrant, and horizontal in the north and south quadrant. So the system indeed interact in both directions, leading to the 2D entanglement we have seen in the previous section regardless of the direction of bipartition. Furthermore, we show in this section that the layout of the 1D chain on the 2D lattice induces an antiferromagnetic order of the spin degree of freedom within each quadrant along the direction that there is no interaction.
The expectation value of the spin in a 1D chain has been computed for $q=1$ from combinatorial methods [30, 31], field theoretic approach in the scaling limit [29], and the transfer matrix method [27]. All three approaches lead to the same scaling behavior as $r^{-1/2}$, where $r$ is the distance from the closer boundary, with up spin on the first half, and down spin on the second. When twisted and embedded on the 2D lattice, an induced antiferromagnetic order appears with modulated amplitude that decays from the boundary to the center, as shown in 2 (b).
As was shown in Ref. [27], due to the opposite boundary conditions, the $q>1$ ordered phase forms a domain wall between the two regions with positive and negative values of the order parameter. The correlation length in this case is reflected in the domain wall thickness $\xi$ by
[…]
which can be derived from the Ginzburg-Landau theory [27]. The length scale depends on the deformation parameter as $\xi\propto 1/\log q$, which was obtained from numerical results by a matrix product state calculation [27]. Fig. 2 (a) depicts the twisting induced antiferromagnetic order in the direction orthogonal to 1D chain within each quadrant of the 2D lattice for $\xi=2N$.
V TN
Figure 3: (a) TN representation of the 2D GS of the twisted Motzkin chain, where the contractions in the vertical direction are represented by placing the tensor cubes adjacent to each other, and in the horizontal planes, contraction only happens between neighbors along the twisted chain direction. (b) TN representation of the 2D GS of the coupled chain model, where contractions in each horizontal plane happen between neighboring tensor cubes in both directions.